Jean paints a house in 4 days. Brett can paint the same house in 6 days and Joe can paint the same house in 2 days. How much time will it take to paint the house together?

Answer by rwm(914)

x/4+x/6+x/2=1
12x/4+12x/6+12x/2=12
3x+2x+6x=12
11x=12
x=12/11 days